# Critical Quantum Sensing: a tutorial on parameter estimation near quantum phase transitions

### v1

@@ IV.6 Precision limits for systems where the total number of excitations is not fixed
This reformulation extends the concepts of the standard quantum limit and the Heisenberg limit to systems where the total number of excitations is not fixed, for instance, in photonic systems. In such cases, $N$ is effectively replaced by $\langle\hat{n}\rangle$, providing a consistent framework for characterizing the ultimate precision limits of these systems.
@@ IV.7 Quantum-enhanced measurements
We can therefore see how the use of quantum coherence and entanglement can be used as a resource for sensing to go beyond the typical precision limits. This is essentially achieved by exploiting the fact that entangled systems collectively respond to small variations in parameters of interest. A potential shortcoming from a practical perspective is the prohibitively large time and resources required to generate squeezed and maximally entangled states [140]. Furthermore, such states are intrinsically sensitive to noise and decoherence effects which ultimately serve to degrade performance below the coveted Heisenberg limit. This motivates the search for alternative platforms which may achieve the Heisenberg limit, without suffering from these prohibitive shortcomings. Critical systems emerge as natural candidates in this regard: many critical systems are “passive” in the sense that, a many-body system close to criticality can be tuned to its critical point and allowed to respond to a small parameter change. In practice, this is can be implemented in condensed matter systems.
@@ V Minimal critical model: the quantum harmonic oscillator
V.1 Harmonic oscillator oscillator: closing the energy gap
@@
where $\omega(g)=\omega\sqrt{1-g/\omega}$ is the tunable frequency, and the parameter $g$ controls the oscillator’s behavior: increasing $g$ opens (unsqueezes) the oscillator, while decreasing it closes (squeezes) the harmonic oscillator (see Fig. 5 for a schmatic illustration). Clearly, when $g=\omega$, the oscillator flattens, and the energy gap vanishes, mimicking the critical point of a quantum phase transition. To describe this in terms of quantum fields, we express the abstract position and momentum operators as
@@
where $\hat{a}$ and $\hat{a}^{\dagger}$ are the annihilation and creation operators, respectively. Substituting these into the Hamiltonian, we obtain
@@
where $\xi=\frac{1}{4}\log\left[1-g/\omega\right]$ is the squeezing parameter. The transformed Hamiltonian, expressed in terms of a new set of annihilation and creation operators $\hat{c}$ and $\hat{c}^{\dagger}$, becomes
@@
Figure 5: Schematic illustrating the connection between gap closing and quantum correlations. Panels (a)–(c) show the progressive closing of the energy gap in a harmonic oscillator, corresponding to the flattening of its potential (d)–(f). This deformation of the potential strongly modifies the ground state. As the oscillator opens up, the state becomes increasingly squeezed (g)-(i). At the critical point, where the harmonic oscillator is maximally flattened, the ground state exhibits extreme squeezing and heightened sensitivity to the oscillator frequency. This mechanism provides an intuitive picture for the enhanced sensitivity underlying critical quantum metrology.
@@ V.2 Quantum Fisher information for an opening harmonic oscillator
The squeezing associated with opening or closing the energy gap already hints at the possibility of reducing quantum noise and thus improving measurement precision, analogous to squeezing noise in Ramsey interferometry. To explore this further, let us calculate the QFI, treating $\omega$ as the unknown parameter for the ground state of the modified harmonic oscillator described earlier. Plugging the ground state from Eq. (50) with $n=0$ into the expression for the QFI (36), we obtain
@@
This result explodes at the critical point $g\sim\omega$ where the harmonic oscillator frequency vanishes as $\omega(g)\to 0$, signaling an infinite QFI. However, this form of QFI is unconventional; notably, it appears independent of time and the number of excitations. Does this imply that the bounds derived in earlier sections do not apply to critical metrology?
@@
$\langle\hat{a}^{\dagger}\hat{a}\rangle=\sinh^{2}\xi\xrightarrow[]{g\rightarrow\omega}\frac{1}{4}\frac{1}{\sqrt{1-g/\omega}}\;.$ (58)
@@
$\mathcal{I}_{\omega}\approx\frac{4g^{2}}{\omega^{4}}\langle\hat{a}^{\dagger}\hat{a}\rangle^{4}\;,$ (59)
@@
$t<\frac{2g}{\omega^{2}}\langle\hat{a}^{\dagger}\hat{a}\rangle\;$. (60)
@@ V.3 Taking time into consideration
where we enforce $\gamma\ll 1$ to satisfy the adiabatic condition. Here, the time $T$ is not a freely tunable parameter as in the case of free evolution; rather, it is a function of the number of excitations $\langle\hat{n}\rangle$, which grows as one approaches the critical point. This connection between time and excitations is a hallmark of critical systems and plays a central role in critical quantum metrology. Near a quantum phase transition, the energy gap closes in the thermodynamic limit, leading to critical slowing down—the phenomenon where the system’s intrinsic timescales diverge as it approaches criticality. As a result, any adiabatic process that attempts to traverse the critical region becomes prohibitively slow, requiring increasingly long times to maintain adiabaticity and avoid diabatic transitions. This divergence is reflected in the scaling of the relaxation time, which typically follows a power-law behavior with the distance to the critical point.
@@
which exhibits a quadratic scaling in both $\langle\hat{n}\rangle$ and TT—commonly referred to as Heisenberg scaling. However, this comes with an important caveat: the prefactor is suppressed by $8\gamma^{2}$, making the achievable precision substantially worse than the Heisenberg limit. In fact, the Fisher information surpasses the standard quantum limit only when
@@
Figure 6: Operating mechanism of critical metrology on the example of an opening harmonic oscillator. Approaching the critical point $g=\omega$ opens (or squeezes) the harmonic oscillator which squeezes the ground state as visible in (a)-(d) where we depict the Husimi function of the ground state. This squeezing leads to the increased signal (here related to $\langle\hat{n}\rangle$) but also to increased fluctuations $(\Delta\hat{n})^{2}$ (e). Close to the critical point. The signal, however, grows much faster than the fluctuations. This leads to an overall enhancement of the SNR (and the QFI). Ideally, one would want to increase the signal and reduce the noise at the same time.
@@ V.4 Optimal observable
So far we have considered only the QFI. Let us now analyze the SNR in the harmonic oscillator model. A natural choice of observables includes the excitation number, $\hat{n}=\hat{a}^{\dagger}\hat{a}$, or the squeezing, represented by the width of the harmonic oscillator, $(\hat{a}+\hat{a}^{\dagger})^{2}$. For the excitation number $\hat{n}$, the fluctuations are given by
@@
where $\langle\hat{n}\rangle=\sinh^{2}\xi$ is the average number of excitations. Substituting the value of $\langle\hat{n}\rangle$ from Eq. (58), we calculate the SNR for measuring the average excitation number
@@
Second, while critical systems are known to amplify noise (as seen in the increased fluctuations of observables), they also amplify the signal even more strongly, leading to significant improvements in metrological performance (see Fig. 6). This resonates with the broader observation that the precision of critical metrology improves with system size, a trend we have previously discussed. However, an ideal scenario would involve increasing the signal while simultaneously minimizing—or at least maintaining—the noise. Striking this balance represents a compelling challenge and an exciting avenue for future research in critical metrology. Later in Sec. X, we will revisit this concept and explore strategies for optimizing the SNR in critical systems.
V.5 Finite size system phase transition
The energy gap can only fully close in the thermodynamic limit, where the system size becomes infinite, a condition that cannot be realized in practical, finite-size systems. As a result, the energy gap remains slightly open, leading to deviations from the idealized harmonic oscillator picture—and Gaussianity—often used to describe the system near criticality. To account for these finite-size effects and the resulting modifications to the harmonic behavior, one can introduce a quartic correction to the harmonic oscillator model. This correction takes the form of an anharmonic term in the Hamiltonian:
@@ V.6 Lessons from the harmonic oscillator and the road ahead
VI Quantum phase transitions, criticality, and critical metrology: elementary Landau Zener model
@@
In this section, we introduce quantum phase transitions through the Landau-Zener model, which serves as a minimal, analytically tractable example capturing the essential features of critical behavior in a driven two-level system. In the subsequent section, we then move on to the transverse field Ising model—a paradigmatic system exhibiting a genuine quantum phase transition in the thermodynamic limit. These models allow us to explore how criticality emerges in quantum systems and how the associated nonclassical features, such as diverging susceptibilities and long-range correlations, can be harnessed to enhance metrological precision
@@ VI.1 Landau-Zener model
Far from the avoided crossing ($|g|\gg\omega$), the transverse field dominates and the eigenstates align with the eigenstates of $\hat{\sigma}_{x}$. In contrast, near $g=0$, the $\hat{\sigma}_{z}$ term becomes dominant, and the system closely approximates the bare $|\!\downarrow\rangle$ and $|\!\uparrow\rangle$ basis. As the control field $g$ is varied adiabatically from large negative to large positive values, the ground state evolves smoothly but significantly, interpolating between two qualitatively distinct regimes. This continuous yet abrupt transformation of the ground state is characteristic of a quantum phase transition. The phases in this context correspond to regions where either the $\hat{\sigma}_{x}$ or $\hat{\sigma}_{z}$ term dominates, while the crossover near $g=0$ serves as an analogue of a quantum critical point. Importantly, the sharpness of this transition depends on $\omega$. A small $\omega$ results in a narrow avoided crossing and a more abrupt change in the ground state, closely resembling a true quantum phase transition [see Figs. 7(a) and 7(b)]. Conversely, larger $\omega$ values smooth out the transition and suppress critical-like features. Despite its simplicity, the LZ model thus encapsulates many of the qualitative hallmarks of quantum criticality, making it a powerful pedagogical tool and a stepping stone toward understanding more complex models, such as the transverse field Ising model.
Figure 7: Properties of the LZ model. (a) Energy eigenvalues of the LZ model. (b) Energy gap $\Delta=\sqrt{\omega^{2}+g^{2}}$ of the LZ model. (c) QFI for estimating $\omega$. (d) Geometric change of the angle $\phi$. (e) Rate of change of angle $\phi$ with respect to $g$ showing maximal rate of change when $\omega=g$ gap closes. (f) Rate of change of angle $\phi$ with respect to $\omega$ illustrating why the QFI does not peak when gap closes $\omega=g=0$ but at some other value: because this corresponds to the path of maximal curvature of the eigenstates. The black cross in (c) and (f) are to emphasize this point and represent the peak QFI and peak $\partial_{\omega}\phi$ respectively.
@@ VI.2 Ground state sensitivity
where $\Delta=\sqrt{g^{2}+\omega^{2}}$ is the instantaneous energy gap. As $\omega\to 0$, the gap closes at $g=0$, leading to a divergence in the QFI. This phenomenon is rooted in the general expression for the QFI in terms of the eigenstates and eigenvalues of the Hamiltonian:
@@ VI.6 Thermal fluctuations
Figure 8: Single-parameter estimation at finite temperature for the LZ model. (a) Contribution of the thermal populations to probe parameter sensitivity $\mathcal{I}_{\text{pop}}$. (b) Contribution of the temperature-dependent weighting factor $\mathcal{W}(T)$ and groundstate sensitivity $\mathcal{I}_{\text{GS}}$. (c) Total finite-temperature QFI for estimating $\omega$. (d) Comparison of the QFI for several values of $g$. The frequency is fixed $\omega=1$ in all panels.
@@
and $\hat{H}|e_{n}\rangle=E_{n}|e_{n}\rangle$ defines the system’s eigenbasis and we use natural units with $k_{B}=1$ throughout. At nonzero temperature, information about the parameter $\omega$ can be imprinted both in the thermal populations $e_{n}$ and in the structure of the eigenstates $|e_{n}\rangle$.
@@
These features are visualized in Fig. 8. Panel (a) shows the contribution of the population dependant term of the total QFI, $\mathcal{I}_{\text{pop}}$, reflecting the degree of information now imprinted on the thermal populations. This quantity peaks at intermediate temperatures $T\sim\Delta$ and small $g$, where both population imbalance and coherent transitions are significant. The maximum occurs at $g=0$ where the crossing occurs and the population change is maximal. Notably, this prefactor vanishes in the high-temperature limit, where thermal populations flatten and coherence is suppressed. Panel (b) illustrates the impact the thermal weighting factor $\mathcal{W}(T)$ has on the groundstate sensitivity $\mathcal{I}_{\text{GS}}$. In contrast with population dependent thermal contribution, here the sensitivity peaks around $\omega=g$ in line with the zero-temperature limit. The sensitivity deteriorates from its maximum as the thermal fluctuations being to dominate.
In panel (c) the full finite-temperature QFI $\mathcal{I}^{\beta}_{\omega}$ is displayed, revealing a broad sensitivity maximum at $g\approx 0$ and $T\sim\Delta$, precisely where thermal and quantum effects optimally combine. This highlights that temperature does not merely degrade sensitivity; it can in fact enable new metrological channels if harnessed appropriately. Panel (d) further emphasizes this by plotting the temperature dependence of QFI for three values of $g$: 0.1 (black solid), 0.5 (orange dotted), and 1 (magenta dashed). In all cases, QFI approaches the groundstate value as $T\to 0$, then exhibits an enhancement in the intermediate-temperature regime before decaying to zero at high temperatures.
Taken together, these results demonstrate that optimal sensitivity need not occur at zero temperature. Instead, a finely tuned balance between quantum coherence and thermal activation can yield superior precision. This reveals a key insight in critical metrology: thermal fluctuations, typically considered detrimental, can under specific conditions act as a resource—particularly near critical points, where energy gaps are small and sensitivity is high.

### v2

@@ IV.6 Precision limits for systems where the total number of excitations is not fixed
This reformulation extends the concepts of the standard quantum limit and the Heisenberg limit to systems where the total number of excitations is not fixed, as photonic systems. In such cases, $N$ is effectively replaced by $\langle\hat{n}\rangle$, providing a consistent framework for characterizing the ultimate precision limits of these systems.
@@ IV.7 Quantum-enhanced measurements
We can therefore see how quantum coherence and entanglement can be used as a resource for sensing to go beyond the typical precision limits. This is essentially achieved by exploiting the fact that entangled systems collectively respond to small variations in parameters of interest. A potential shortcoming from a practical perspective is the prohibitively large time and resources required to generate squeezed and maximally entangled states [153]. Furthermore, such states are intrinsically sensitive to noise and decoherence, which ultimately degrade their performance below the Heisenberg limit. This motivates the search for alternative platforms which may achieve the Heisenberg limit, without suffering from these prohibitive shortcomings. Critical systems emerge as natural candidates in this regard: many critical systems are “passive” in the sense that a many-body system close to criticality can be tuned to its critical point and allowed to respond to a small parameter change. In practice, this is can be implemented in condensed matter systems.
@@ V Minimal critical model: the quantum harmonic oscillator
V.1 Harmonic oscillator: closing the energy gap
@@
where $\omega(g)=\omega\sqrt{1-g/\omega}$ is the tunable frequency, and the parameter $g$ controls the oscillator’s behavior: increasing $g$ opens (unsqueezes) the oscillator, while decreasing it closes (squeezes) the harmonic oscillator (see Fig. 5 for a schematic illustration). Clearly, when $g=\omega$ the oscillator flattens, and the energy gap vanishes, mimicking the critical point of a quantum phase transition. To describe this in terms of quantum fields, we express the abstract position and momentum operators as
@@
where $\hat{a}$ and $\hat{a}^{\dagger}$ are the annihilation and creation operators, respectively. Substituting these into Eq. (47), we obtain
@@
where $\xi=\left[\log\left(1-g/\omega\right)\right]/4$ is the squeezing parameter. The transformed Hamiltonian, expressed in terms of a new set of annihilation and creation operators $\hat{c}$ and $\hat{c}^{\dagger}$, becomes
@@
Figure 5: Schematic illustration of the connection between energy-gap closing and quantum correlations. Panels (a)–(c) show the progressive closing of the energy gap in a harmonic oscillator, corresponding to the flattening of its potential (d)–(f). This deformation of the potential strongly modifies the ground state. As the oscillator opens up, the state becomes increasingly squeezed (g)-(i). At the critical point, where the harmonic oscillator is maximally flattened, the ground state exhibits extreme squeezing and heightened sensitivity to the oscillator frequency. This mechanism provides an intuitive picture for the enhanced sensitivity underlying critical quantum metrology.
@@ V.2 Quantum Fisher information for an opening harmonic oscillator
The squeezing associated with opening or closing of the energy gap already hints at the possibility of reducing quantum noise and thus improving measurement precision, in analogy to noise squeezing in Ramsey interferometry. To further explore this, let us calculate the QFI for the ground state of the modified harmonic oscillator previously described, treating $\omega$ as the unknown parameter. Plugging the ground state from Eq. (50) with $n=0$ into the expression for the QFI (36), we obtain
@@
This result explodes at the critical point $g\sim\omega$ where the harmonic oscillator frequency vanishes as $\omega(g)\to 0$, signaling an infinite QFI. However, this form of the QFI is unconventional; notably, it appears independent of time and the number of excitations. Does this imply that the bounds derived in earlier sections do not apply to critical metrology?
@@
$\langle\hat{n}\rangle=\langle\hat{a}^{\dagger}\hat{a}\rangle=\sinh^{2}\xi\xrightarrow[]{g\rightarrow\omega}\frac{1}{4}\frac{1}{\sqrt{1-g/\omega}}\;.$ (58)
@@
$\mathcal{I}_{\omega}\approx\frac{4g^{2}}{\omega^{4}}\langle\hat{n}\rangle^{4}\;,$ (59)
@@
$t<\frac{2g}{\omega^{2}}\langle\hat{n}\rangle\;$. (60)
@@ V.3 Taking time into consideration
where the condition $\gamma\ll 1$ is imposed to ensure the adiabatic condition. Here, the time $T$ is not a freely tunable parameter as in the case of free evolution; rather, it is a function of the number of excitations $\langle\hat{n}\rangle$, which grows as one approaches the critical point. This connection between time and excitations is a hallmark of critical systems and plays a central role in critical quantum metrology. Near a quantum phase transition, the energy gap closes in the thermodynamic limit, leading to critical slowing down—the phenomenon where the system’s intrinsic timescales diverge as it approaches criticality. As a result, any adiabatic process that attempts to traverse the critical region becomes prohibitively slow, requiring increasingly long times to maintain adiabaticity and avoid diabatic transitions. This divergence is reflected in the scaling of the relaxation time, which typically follows a power-law behavior with the distance to the critical point.
@@
which exhibits a quadratic scaling in both $\langle\hat{n}\rangle$ and TT—commonly referred to as Heisenberg scaling. However, this comes with an important caveat: the prefactor is suppressed by $8\gamma^{2}$, making the achievable precision substantially worse than the Heisenberg limit. In fact, the QFI surpasses the standard quantum limit only when
@@
Figure 6: Operating mechanism of critical metrology illustrated with an opening harmonic oscillator. Approaching the critical point $g=\omega$ opens (or squeezes) the harmonic oscillator which in turn squeezes the ground state, as shown in panels (a)–(d), where we display the Husimi Q-function of the ground state. This squeezing increases the signal (here related to $\langle\hat{n}\rangle$) but also enhances fluctuations $(\Delta\hat{n})^{2}$, as shown in (e). However, near the critical point, the signal grows much faster than the fluctuations, leading to an overall enhancement of the SNR and the QFI. Ideally, one would aim to enhance the signal while simultaneously suppressing the noise.
@@ V.4 Optimal observable
So far we have considered only the QFI. Let us now analyze the SNR in the harmonic oscillator model. A natural choice of observables includes the excitation number $\hat{n}=\hat{a}^{\dagger}\hat{a}$, or the squeezing represented by the width of the harmonic oscillator, $(\hat{a}+\hat{a}^{\dagger})^{2}$. For the excitation number $\hat{n}$, the fluctuations are given by
@@
where $\langle\hat{n}\rangle=\sinh^{2}\xi$ is the average number of excitations. Substituting the value of $\langle\hat{n}\rangle$ from Eq. (58), we calculate the SNR for measuring the average excitation number:
@@
Second, while critical systems are known to amplify noise (as seen in the increased fluctuations of observables), they also amplify the signal even more strongly, leading to significant improvements in metrological performance (see Fig. 6). This resonates with the broader observation that the precision of critical metrology improves with system size, a trend we have previously discussed. However, an ideal scenario would involve increasing the signal while simultaneously minimizing—or at least maintaining—the noise. Striking this balance represents a compelling challenge and an exciting avenue for future research in critical metrology. Later in Section X, we will revisit this concept and explore strategies for optimizing the SNR in critical systems.
V.5 Finite-size system phase transition
The energy gap can only fully close in the thermodynamic limit where the system size becomes infinite, a condition that cannot be realized in practical, finite-size systems. As a result, the energy gap remains slightly open, leading to deviations from the idealized harmonic oscillator picture—and Gaussianity—often used to describe the system near criticality. To account for these finite-size effects and the resulting modifications to the harmonic behavior, one can introduce a quartic correction to the harmonic oscillator model. This correction takes the form of an anharmonic term in the Hamiltonian:
@@ V.6 Lessons from the harmonic oscillator and the road ahead
VI Quantum phase transitions, criticality, and critical metrology: elementary Landau-Zener model
@@
In this section, we introduce quantum phase transitions through the Landau-Zener model, which serves as a minimal, analytically tractable example capturing the essential features of critical behavior in a driven two-level system. In the subsequent section, we then move on to the transverse-field Ising model—a paradigmatic system exhibiting a genuine quantum phase transition in the thermodynamic limit. These models allow us to explore how criticality emerges in quantum systems and how the associated nonclassical features, such as diverging susceptibilities and long-range correlations, can be harnessed to enhance metrological precision
@@ VI.1 Landau-Zener model
Far from the avoided crossing ($|g|\gg\omega$), the transverse field dominates and the eigenstates align with the eigenstates of $\hat{\sigma}_{x}$. In contrast, near $g=0$, the $\hat{\sigma}_{z}$ term becomes dominant, and the system closely approximates the bare $|\!\downarrow\rangle$ and $|\!\uparrow\rangle$ basis. As the control field $g$ is varied adiabatically from large negative to large positive values, the ground state evolves smoothly but significantly, interpolating between two qualitatively distinct regimes. This continuous yet abrupt transformation of the ground state is characteristic of a quantum phase transition. The phases in this context correspond to regions where either the $\hat{\sigma}_{x}$ or $\hat{\sigma}_{z}$ term dominates, while the crossover near $g=0$ serves as an analogue of a quantum critical point. Importantly, the sharpness of this transition depends on $\omega$. A small $\omega$ results in a narrow avoided crossing and a more abrupt change in the ground state, closely resembling a true quantum phase transition [see Figs. 7(a) and 7(b)]. Conversely, larger $\omega$ values smooth out the transition and suppress critical-like features. Despite its simplicity, the LZ model thus encapsulates many of the qualitative hallmarks of quantum criticality, making it a powerful pedagogical tool and a stepping stone toward understanding more complex models, such as the transverse-field Ising model.
Figure 7: Properties of the LZ model. (a) Energy eigenvalues of the LZ model. (b) Energy gap $\Delta=\sqrt{\omega^{2}+g^{2}}$ of the LZ model. (c) QFI for estimating $\omega$. (d) Geometric change of the angle $\phi$. (e) Rate of change of angle $\phi$ with respect to $g$, showing maximal rate of change when $\omega=g$, where the energy gap closes. (f) Rate of change of angle $\phi$ with respect to $\omega$ illustrating why the QFI does not peak at $\omega=g=0$ (when the energy gap closes) but rather at a finite value corresponding to the point of maximal curvature of the eigenstate manifold. The black cross in (c) and (f) are to emphasize this point and represent the peak of the QFI and $\partial_{\omega}\phi$, respectively.
@@ VI.2 Ground state sensitivity
where $\Delta=\sqrt{g^{2}+\omega^{2}}$ is the instantaneous energy gap. As $\omega\to 0$, the gap closes at $g=0$, leading to a divergence in the QFI. This phenomenon is rooted in the general expression for the QFI in terms of $\omega$-dependent eigenstates $|\psi_{n}(\omega)\rangle$ and eigenvalues $E_{n}(\omega)$ of the Hamiltonian $\hat{H}(\omega)$:
@@ VI.6 Thermal fluctuations
Figure 8: Single-parameter estimation at finite temperature for the LZ model. (a) Contribution of the thermal populations to probe parameter sensitivity $\mathcal{I}_{\text{pop}}$. (b) Contribution of the temperature-dependent weighting factor $\mathcal{W}(T)$ and ground state sensitivity $\mathcal{I}_{\text{GS}}$. (c) Total finite-temperature QFI for estimating $\omega$. (d) Comparison of the QFI for several values of $g$. In all panels we set $\omega=1$.
@@
where $\hat{H}|e_{n}\rangle=E_{n}|e_{n}\rangle$ defines the system’s eigenbasis. Throughout the manuscript, we use natural units with $k_{B}=1$. At non-zero temperature $T$, information about the parameter $\omega$ can be imprinted both in the thermal populations $e_{n}$ and in the structure of the eigenstates $|e_{n}\rangle$.
@@
These features are visualized in Fig. 8. Panel (a) shows the contribution of the population dependant term of the total QFI, $\mathcal{I}_{\text{pop}}$, reflecting the degree of information now imprinted on the thermal populations. This quantity peaks at intermediate temperatures $T\sim\Delta$ and small $g$, where both population imbalance and coherent transitions are significant. The maximum occurs at $g=0$ where the crossing occurs and the population change is maximal. Notably, this prefactor vanishes in the high-temperature limit, where thermal populations flatten and coherence is suppressed. Panel (b) illustrates the impact the thermal weighting factor $\mathcal{W}(T)$ has on the ground state sensitivity $\mathcal{I}_{\text{GS}}$. In contrast with the population-dependent thermal contribution, here the sensitivity peaks around $\omega=g$ in line with the zero-temperature limit. The sensitivity deteriorates from its maximum as the thermal fluctuations being to dominate.
In panel (c) the full finite-temperature QFI $\mathcal{I}^{\beta}_{\omega}$ is displayed, revealing a broad sensitivity maximum at $g\approx 0$ and $T\sim\Delta$, precisely where thermal and quantum effects optimally combine. This highlights that temperature does not merely degrade sensitivity; it can in fact enable new metrological channels if harnessed appropriately. Panel (d) further emphasizes this by plotting the temperature dependence of QFI for three values of $g=\{0.1,0.5,1\}$. In all cases, the QFI approaches the ground state value as $T\to 0$, then exhibits an enhancement in the intermediate-temperature regime before decaying to zero at high temperatures.
Taken together, these results demonstrate that optimal sensitivity does not occur at zero temperature. Instead, a finely tuned balance between quantum coherence and thermal activation can yield superior precision. This reveals a key insight in critical metrology: thermal fluctuations, typically considered detrimental, can under specific conditions act as a resource—particularly near critical points, where energy gaps are small and sensitivity is high.